# Rethinking Transformer-Based Blind-Spot Network for Self-Supervised Image Denoising

### v3

@@ Introduction
Recently, self-supervised image denoising (SSID) (Krull, Buchholz, and Jug 2019; Lee, Son, and Lee 2022; Li et al. 2023; Wang et al. 2023b; Jang et al. 2023a; Zhang et al. 2023) has been introduced to circumvent the requirement of the paired datasets. The pioneering work Noise2Void randomly masks some locations of a noisy input and trains the network to reconstruct them from their surrounding ones. In the case of pixel-wise independent noise, the network learns to predict the masked pixels without random noise, i.e., the clean pixels, thereby achieving self-supervised denoising. Blind-spot networks (BSN) (Laine et al. 2019; Wu et al. 2020; Byun, Cha, and Moon 2021) take a step further to implement the mask schema with dedicated designed network architectures that exclude the corresponding input pixel from the receptive field of each output location, which shows superiority in both performance and training efficiency. For spatially correlated noise removal in real-world scenarios, some works (Wu et al. 2020; Zhou et al. 2020) suggest first breaking the noise correlation with pixel-shuffle downsampling (PD), and then denoising with BSN. In particular, asymmetric PD strategy (Lee, Son, and Lee 2022; Wang et al. 2023b; Jang et al. 2023a) for training and inference has shown a better trade-off between noise removal and detail preserving.
@@
The development of learning-based methods (Zhang et al. 2017) has shown superior performance against traditional patch-based ones (Buades, Coll, and Morel 2005; Dabov et al. 2007) on synthetic Gaussian denoising. More advanced deep neural architectures (Mao, Shen, and Yang 2016; Tai et al. 2017; Liu et al. 2018) are further proposed to improve the denoising ability. NBNet (Cheng et al. 2021) proposes a noise basis network by learning a set of reconstruction basis in the feature space. InvDN (Liu et al. 2021a) proposes a lightweight denoising network based on normalizing flow architectures. Recently, transformers that were first introduced for sequence modeling in natural language processing (Vaswani et al. 2017) have been successfully applied to vision tasks (Dosovitskiy et al. 2020; Liu et al. 2021b). For image denoising, transformers are studied with large-scale image pre-training (Chen et al. 2021) and Swin Transformer architectures (Liang et al. 2021). Restormer (Zamir et al. 2022) and Uformer (Wang et al. 2022a) propose multi-scale hierarchical network designs, which achieve better trade-offs between performance and efficiency. However, there are limited efforts for adapting transformers into self-supervised image denoising (Wang et al. 2023b; Papkov and Chizhov 2023) due to the blind-spot requirement.
@@
As shown in Fig. 2, TBSN follows dilated BSN (Wu et al. 2020) to apply 3$\times$3 centrally masked convolution at the first layer and dilated transformer attention blocks (DTABs) in the remaining layers. The network architecture is U-Net and adopts patch-unshuffle/shuffle (Jang et al. 2023a) based downsampling/upsampling operations to maintain the blind-spot requirement. The building block, i.e., DTAB, is formed with dilated counterparts of grouped channel-wise self-attention (G-CSA), masked window-based self-attention (M-WSA), and feed-forward network (FFN), respectively. Thus, TBSN benefits from both the global interaction of channel attention and the local fitting ability of window attention. We will provide detailed illustrations of the network design in the following subsections.
@@
[…]
@@
Specifically, $\mathbf{M}$ is a two-valued matrix that masks out certain locations according to the relative position of query (at $i$) and key/value (at $j$) tokens. $(x_{i},y_{i})$ and $(x_{j},y_{j})$ are the spatial locations of $i$ and $j$. When $i$ and $j$ are with even distance on both axes, $\mathbf{M}(i,j)\!=\!0$, the attention value is unchanged. Otherwise, $\mathbf{M}(i,j)\!=\!-\infty$ and the attention value becomes 00 after the softmax operation, thereby being masked out. Inspired by relative position embedding (Liu et al. 2021b), $\mathbf{M}\in\mathbb{R}^{M^{2}\times M^{2}}$ can be calculated from a smaller-sized binary matrix $\hat{\mathbf{M}}\in\mathbb{R}^{(2M-1)\times(2M-1)}$ according to the relative position of $i$ and $j$ to improve the efficiency, i.e.,
@@
Supervised (Synthetic pairs) DnCNN 26.25 / 0.599 32.43 / 0.790
@@
Supervised (Real pairs) DnCNN 37.61 / 0.941 38.73 / 0.945
@@
Self-Supervised (Ours) TBSN 37.78 / 0.940
@@
Table 3: Ablation study of dilated transformer attention block (DTAB) on SIDD validation dataset. ’-’ refers to plain dilated convolutions. (1$\sim$3) indicates our dilated M-WSA and G-CSA are essential components of DTAB. (4$\sim$8) shows they are not replaceable with analogous operators.

### v4

@@ Introduction
Recently, self-supervised image denoising (SSID) (Krull, Buchholz, and Jug 2019; Lee, Son, and Lee 2022; Li et al. 2023; Wang et al. 2023b; Jang et al. 2023a) has been introduced to circumvent the requirement of the paired datasets. The pioneering work Noise2Void randomly masks some locations of a noisy input and trains the network to reconstruct them from their surrounding ones. In the case of pixel-wise independent noise, the network learns to predict the masked pixels without random noise, i.e., the clean pixels, thereby achieving self-supervised denoising. Blind-spot networks (BSN) (Laine et al. 2019; Wu et al. 2020; Byun, Cha, and Moon 2021) take a step further to implement the mask schema with dedicated designed network architectures that exclude the corresponding input pixel from the receptive field of each output location, which shows superiority in both performance and training efficiency. For spatially correlated noise removal in real-world scenarios, some works (Wu et al. 2020; Zhou et al. 2020) suggest first breaking the noise correlation with pixel-shuffle downsampling (PD), and then denoising with BSN. In particular, asymmetric PD strategy (Lee, Son, and Lee 2022; Wang et al. 2023b; Jang et al. 2023a) for training and inference has shown a better trade-off between noise removal and detail preserving.
@@
The development of learning-based methods (Zhang et al. 2017) has shown superior performance against traditional patch-based ones (Buades, Coll, and Morel 2005; Dabov et al. 2007) on synthetic Gaussian denoising. More advanced deep neural architectures (Mao, Shen, and Yang 2016; Tai et al. 2017; Liu et al. 2018) are further proposed to improve the denoising ability. NBNet (Cheng et al. 2021) proposes a noise basis network by learning a set of reconstruction basis in the feature space. InvDN (Liu et al. 2021a) proposes a lightweight denoising network based on normalizing flow architectures. Recently, transformers that were first introduced for sequence modeling in natural language processing (Vaswani et al. 2017) have been successfully applied to vision tasks (Dosovitskiy et al. 2020; Liu et al. 2021b). For image denoising, transformers are studied with large-scale image pre-training (Chen et al. 2021) and Swin Transformer architectures (Liang et al. 2021). Restormer (Zamir et al. 2022) and Uformer (Wang et al. 2022) propose multi-scale hierarchical network designs, which achieve better trade-offs between performance and efficiency. However, there are limited efforts for adapting transformers into self-supervised image denoising (Wang et al. 2023b; Papkov and Chizhov 2023) due to the blind-spot requirement.
@@
As shown in Fig. 2, TBSN follows dilated BSN (Wu et al. 2020) to apply $\times3$ centrally masked convolution at the first layer and dilated transformer attention blocks (DTABs) in the remaining layers. The network architecture is U-Net and adopts patch-unshuffle/shuffle (Jang et al. 2023a) based downsampling/upsampling operations to maintain the blind-spot requirement. The building block, i.e., DTAB, is formed with dilated counterparts of grouped channel-wise self-attention (G-CSA), masked window-based self-attention (M-WSA), and feed-forward network (FFN), respectively. Thus, TBSN benefits from both the global interaction of channel attention and the local fitting ability of window attention. We will provide detailed illustrations of the network design in the following subsections.
@@
$\text{G-CSA}(\mathbf{X})=\text{Concat}(\mathbf{X}_{1}\ast\phi(\mathbf{X}_{1}),\cdots,\mathbf{X}_{G}\ast\phi(\mathbf{X}_{G})).$ (2)
@@
Specifically, $\mathbf{M}$ is a two-valued matrix that masks out certain locations according to the relative position of query (at $i$) and key/value (at $j$) tokens. $(x_{i},y_{i})$ and $(x_{j},y_{j})$ are the spatial locations of $i$ and $j$. When $i$ and $j$ are with even distance on both axes, $\mathbf{M}(i,j)\!=\!0$, the attention value is unchanged. Otherwise, $\mathbf{M}(i,j)\!=\!$-$\infty$ and the attention value becomes 0 after the softmax operation, thereby being masked out. Inspired by relative position embedding (Liu et al. 2021b), $\mathbf{M}\in\mathbb{R}^{M^{2}\times M^{2}}$ can be calculated from a smaller-sized binary matrix $\hat{\mathbf{M}}\in\mathbb{R}^{(2M-1)\times(2M-1)}$ according to the relative position of $i$ and $j$ to improve the efficiency, i.e.,
@@
Supervised
(Synthetic pairs)
DnCNN 26.25 / 0.599 32.43 / 0.790
@@
Supervised
(Real pairs)
DnCNN 37.61 / 0.941 38.73 / 0.945
@@
Self-Supervised
TBSN 37.78 / 0.940